# A Critical Assessment of Visual Sound Source Localization Models Including Negative Audio

### v1

@@ Abstract
The task of Visual Sound Source Localization (VSSL) involves identifying the location of sound sources in visual scenes, integrating audio-visual data for enhanced scene understanding. Despite advancements in state-of-the-art (SOTA) models, we observe three critical flaws: i) The evaluation of the models is mainly focused in sounds produced by objects that are visible in the image, ii) The evaluation often assumes a prior knowledge of the size of the sounding object, and iii) No universal threshold for localization in real-world scenarios is established, as previous approaches only consider positive examples without accounting for both positive and negative cases. In this paper, we introduce a novel test set and metrics designed to complete the current standard evaluation of VSSL models by testing them in scenarios where none of the objects in the image corresponds to the audio input, i.e. a negative audio. We consider three types of negative audio: silence, noise and offscreen. Our analysis reveals that numerous SOTA models fail to appropriately adjust their predictions based on audio input, suggesting that these models may not be leveraging audio information as intended. Additionally, we provide a comprehensive analysis of the range of maximum values in the estimated audio-visual similarity maps, in both positive and negative audio cases, and show that most of the models are not discriminative enough, making them unfit to choose a universal threshold appropriate to perform sound localization without any a priori information of the sounding object, that is, object size and visibility.
@@
Pioneering works [4, 6, 10] addressed VSSL by localizing the sound sources in a given image using classical machine learning techniques. Since then, many learning-based models have been proposed. Most of the works in the literature are assessed in scenarios with a single sounding object that is visible in the input image, e.g. [1, 2, 3, 18, 12, 15, 16, 21, 22, 23, 24]. This is not true in practical scenarios where it is common to encounter silent objects, sounds emitted by non visible objects (i.e. offscreen sounds) and different sources sounding at the same time (i.e. a mixture of sounds).
@@
We propose a more comprehensive evaluation of VSSL models, that does not assume any prior information about the sounding objects and that includes not only positive audio examples but also different types of negative audio, namely: silence, noise and offscreen sounds. Consequently, we propose new metrics that assess the performance both in the negative cases and globally in positive and negative cases. By examining the overlap of model prediction maps for both positive and negative cases, we identify which models are more discriminative and can be better adapted to work in real-world applications via an absolute threshold. Our contributions are the following: i) An extended test set for evaluating VSSL models that includes both positive and negative audio examples; ii) New metrics to evaluate the performance of VSSL models on negative audio alone and to assess the global performance both on positive and negative audio; iii) A comprehensive analysis of VSSL models on their discriminatory power between sounding and non sounding objects, for which we identify model-specific universal thresholds to perform VSSL without any prior information of the sounding object related to its size and its visibility. Our code is open source and available in https://github.com/xavijuanola/vssl_eval.
@@ II Method
Extended test datasets. To assess the performance of VSSL models both with positive and negative audio we extend two existing testing datasets: the VGG-SS test set [3] and the recently proposed synthetic IS3 dataset [22]. For negative cases, the models’ localization map (i.e. a binarized map derived from likelihood values assigned to image pixels indicating the presence of a sounding object) should be empty (i.e. low values). We propose three types of negative audio: silence, noise and offscreen sounds. We pair every image in the test set with its positive audio and the three types of negative audio. VGG-SS features 221 categories and IS3 130 categories which are a subset from the VGG-SS ones. However, there are many categories highly related semantically: e.g. there are 6 categories related to dog (e.g. ‘dog barking’, ‘dog bow-wow’ or ‘dog howling’), 3 related to train (‘train wheels squealing’, ‘train whistling’ and ‘train horning’), 2 related to woman (‘female singing’ and ‘female speech, woman speaking’), and so on. These categories are impossible to distinguish from each other within the visual modality when we just consider a single image. Motivated by this, we decided to define a set of broad categories for both datasets in order to select a true offscreen audio for a given category that also belongs to a completely different semantic category. The 8 broad semantic categories that we propose are the following: music/musical instruments, human voice, vehicles, devices/machines, animals, weapons, nature ambient sounds and others/miscellaneous. The categories are distinct enough (e.g. animals vs. instruments instead of cats vs. dogs) and thus offscreen sounds are easily distinguishable from onscreen sounds so that a reasonable VSSL model should be able to discriminate them.
In the case of noise, we generate a sequence of random values with normal distribution with zero mean and standard deviation 1. These audios are then normalized with the same functions as the positive ones. For silence, we pair the images with an empty audio. Our extended VGG-SS constitutes a total of 5,537 images, 4.61 hours of positive audio and 13.84 hours of negative audio. For the extended IS3 we have 3,242 distinct images and 6,484 audio files, corresponding to 5.40 hours of positive audio (one audio for each of the two visible sources in each image) and 16.21 hours of negative audio. For a fair comparison, we ensure that all models are tested against the same noise and offscreen audio samples. Given that the randomness in the assignments of offscreen sounds and noise generation can affect our experiments, we repeat them 10 times with different seeds and report the average metrics.
@@ II-3 F and F
| Test set | Model | cIoU | cIoU Adap. | AUC | AUC Adap. | pIA | AUC | pIA | AUC | pIA | AUC | F | F | |||
| VGG-SS | RCGrad [24] | 12.28 | 38.16 | 12.57 | 38.22 | 4.41 | 95.86 | 4.42 | 95.84 | 4.41 | 95.84 | 21.76 | 22.23 | |||
| LVS [3] | 7.88 | 36.25 | 9.48 | 39.02 | 6.05 | 93.81 | 0.60 | 99.36 | 3.47 | 96.38 | 14.60 | 17.26 | ||||
| EZ-VSL [15] | 19.89 | 42.23 | 20.92 | 42.00 | 1.41 | 98.51 | 1.41 | 98.51 | 5.37 | 94.52 | 33.11 | 34.42 | ||||
| FNAC [23] | 21.29 | 45.17 | 22.25 | 42.94 | 5.54 | 94.43 | 5.54 | 94.43 | 7.33 | 92.62 | 34.71 | 35.97 | ||||
| SLAVC [16] | 23.25 | 47.27 | 24.20 | 44.69 | 4.11 | 95.85 | 4.11 | 95.85 | 9.13 | 90.82 | 37.30 | 38.51 | ||||
| SSL-TIE [12] | 22.25 | 48.26 | 22.86 | 45.82 | 0.79 | 99.15 | 0.54 | 99.40 | 3.34 | 96.53 | 36.29 | 37.10 | ||||
| SSL-Align [22] | 16.48 | 48.02 | 17.59 | 44.38 | 2.36 | 97.53 | 0.94 | 98.96 | 2.60 | 97.27 | 28.21 | 29.83 | ||||
| IS3 | RCGrad [24] | 0.10 | 1.07 | 2.54 | 3.57 | 4.49 | 95.81 | 4.49 | 95.77 | 4.47 | 95.83 | 0.20 | 4.95 | |||
| LVS [3] | 8.12 | 39.38 | 9.76 | 41.07 | 4.59 | 95.23 | 0.41 | 99.53 | 1.67 | 98.18 | 15.02 | 17.75 | ||||
| EZ-VSL [15] | 9.50 | 42.08 | 11.01 | 42.66 | 0.99 | 98.95 | 0.99 | 98.95 | 2.66 | 97.21 | 17.31 | 19.79 | ||||
| FNAC [23] | 17.75 | 49.01 | 19.01 | 45.71 | 5.94 | 94.02 | 5.94 | 94.02 | 5.06 | 94.88 | 29.88 | 31.64 | ||||
| SLAVC [16] | 16.29 | 45.05 | 17.64 | 42.96 | 4.37 | 95.94 | 4.37 | 95.94 | 8.34 | 91.61 | 27.79 | 29.72 | ||||
| SSL-TIE [12] | 18.89 | 49.29 | 19.83 | 46.75 | 0.48 | 99.43 | 0.27 | 99.65 | 3.13 | 96.76 | 31.70 | 33.02 | ||||
| SSL-Align [22] | 28.50 | 63.10 | 29.19 | 52.40 | 4.51 | 95.32 | 1.73 | 98.11 | 2.01 | 97.88 | 44.08 | 44.89 | ||||
@@
As it can be observed in Fig. 1, most models present a strong overlap in the range of maximum similarities for the cases of positive, silence and offscreen audio, thus making it difficult to choose a universal threshold in which the negative audios will not be activated in the localization map. SSL-TIE and SSL-Align are the ones that present a higher discriminatory power. SSL-TIE is the most discriminative model in VGG-SS. In IS3, SSL-TIE is good at discriminating silence and noise but less good in case of offscreen. On the other hand, SSL-Align is good at discriminating noise and offscreen in IS3 but less good at silence. In light of the results in Fig. 1, we decide to set the universal localization threshold for each model as the the 75th percentile, or 3rd quartile, of the maximum similarity values of the three types of negative audio together. This value is computed in the test set of VGG-SS and is shown in Fig. 1 (both in numbers and as a horizontal line). We use this threshold to binarize the audio-visual similarity maps and thus obtain the localization map used in all metrics. The threshold is computed in VGG-SS and used in the test sets of both VGG-SS and IS3; in this latter case with the aim of assessing if it generalizes. The threshold is computed differently for RCGrad model, as it uses gradCAM, with output values normalized in the interval . In this case, we use a threshold of .
@@ III Experiments
We conducted our experiments on the extended test sets of VGG-SS [3] and IS3 [22] datasets. Table I contains the results on the positive and each of the three negative audio inputs as well as the global metrics for both types of inputs, the best results are in bold and the second best underlined. The metrics are computed using two different thresholds: our proposed universal threshold (introduced in Section II), and the adaptive threshold proposed in [22] (metrics corresponding to this case include the word Adap.). In the case of positive audio input and IS3 test set, the first and second best models are SSL-Aling and SSL-TIE, respectively, regardless of the threshold used to get the localization maps. In contrast, in VGG-SS, the best model with the universal threshold is SLAVC, and SSL-TIE the second best. With the adaptive threshold, i.e. the one that uses the prior knowledge of the sounding object’s size, the best results are achieved by SSL-TIE. For negative cases, SSL-TIE presents the best results in silence and noise in both datasets. In offscreen sounds, SSL-Align is the best one in VGG-SS and second best in IS3. LVS is the best in offscreen in IS3, although it has a poor result in the positive audios and this is reflected in the global metrics. The global performance for both positive and negative audio is measured by F and F. We observe that best overall model in VGG-SS is SLAVC and in IS3 is SSL-Align, while SSL-TIE is the second best in both datasets. The results in Table I are consistent with the analysis shown in Figure 1, where the best models according to the global metrics are the ones with less overlap in the IQR of maximum similarity values in the positive and negative cases.
@@
A qualitative example is shown in Figure 2, where each row represents a different model, and each column corresponds to different audio inputs paired with the same image. The example is from IS3 and there are two different positive audio files (‘man’ and ‘ukulele’); we also consider the three types of negative audio samples. To show the effect of the universal threshold we set to zero the audio-visual similarity values below that threshold. The resulting values are then normalized to the range and they are combined with the original image. Although SSL-Align is the best overall model in IS3, it fails in this example with the audio corresponding to ‘man’ and ‘silence’. SSL-TIE correctly localizes part of the regions corresponding to both positive audio and produces empty maps for the negative audio samples. Both SLAVC and FNAC produce empty maps for the negatives as well but barely localize the man. In contrast, LVS highlights practically the same region for three different audio inputs: man, ukulele and silence. Finally, RCGrad outputs the same localization regardless of the audio input.
@@
Figure 3 shows the IoU values computed between different pairs of localization maps (positive vs. each type of negatives and negatives vs. negatives). In an effective model, the IoU values for positive vs. negatives should be low, and high for negatives vs. negatives, as the model is expected to produce an empty localization map for negative cases while correctly identifying the sounding region for the positive ones. In VGG-SS, SSL-TIE presents the lowest IoU for positive vs. negatives (and a more pronounced difference with respect to the IoU of negatives vs. negatives). In IS3, according to these same observations, SSL-Align is the best and SSL-TIE is the second best. These results are in accordance with the F values reported in Table I.

### v2

@@ Abstract
The task of Visual Sound Source Localization (VSSL) involves identifying the location of sound sources in visual scenes, integrating audio-visual data for enhanced scene understanding. Despite advancements in state-of-the-art (SOTA) models, we observe three critical flaws: i) The evaluation of the models is mainly focused in sounds produced by objects that are visible in the image, ii) The evaluation often assumes a prior knowledge of the size of the sounding object, and iii) No universal threshold for localization in real-world scenarios is established, as previous approaches only consider positive examples without accounting for both positive and negative cases. In this paper, we introduce extended test set s and new metrics designed to complete the current standard evaluation of VSSL models by testing them in scenarios where none of the objects in the image corresponds to the audio input, i.e. a negative audio. We consider three types of negative audio: silence, noise and offscreen. Our analysis reveals that numerous SOTA models fail to appropriately adjust their predictions based on audio input, suggesting that these models may not be leveraging audio information as intended. Additionally, we provide a comprehensive analysis of the range of maximum values in the estimated audio-visual similarity maps, in both positive and negative audio cases, and show that most of the models are not discriminative enough, making them unfit to choose a universal threshold appropriate to perform sound localization without any a priori information of the sounding object, that is, object size and visibility.
@@
Pioneering works [4, 6, 10] addressed VSSL by localizing the sound sources in a given image using classical machine learning techniques. Since then, many learning-based models have been proposed. Most of the works in the literature are assessed in scenarios with a single sounding object that is visible in the input image, e.g. [1, 2, 3, 18, 12, 15, 16, 21, 22, 23, 24]. This is not true in practical scenarios where it is common to encounter silent objects, sounds emitted by objects out of the image frame (i.e. offscreen sounds) and different sources sounding at the same time (i.e. a mixture of sounds).
@@
We propose a more comprehensive evaluation of VSSL models, that does not assume any prior information about the sounding objects and that includes not only positive audio examples but also different types of negative audio, namely: silence, noise and offscreen sounds. Consequently, we propose new metrics that assess the performance both in the negative cases and globally in positive and negative cases. By examining the overlap of model prediction maps for both positive and negative cases, we identify which models are more discriminative and can be better adapted to work in real-world applications via an absolute threshold. Our contributions are the following: i) Two extended test set s for evaluating VSSL models that includes both positive and negative audio examples; ii) New metrics to evaluate the performance of VSSL models on negative audio alone and to assess the global performance both on positive and negative audio; iii) A comprehensive analysis of VSSL models on their discriminatory power between sounding and non sounding objects, for which we identify model-specific universal thresholds to perform VSSL without any prior information of the sounding object related to its size and its visibility. Our code is open source and available in https://github.com/xavijuanola/vssl_eval.
@@ II Method
Extended test datasets. To assess the performance of VSSL models both with positive and negative audio we extend two existing testing datasets: the VGG-SS test set [3] and the recently proposed synthetic IS3 dataset [22]. For that, we include three types of negative audio: silence, noise and offscreen sounds. We pair every image in the test set with its positive audio (the original one) and the three types of negative audio that we add. For negative cases, the models’ localization map (i.e. a binarized map derived from likelihood values assigned to image pixels indicating the presence of a sounding object) should be empty (i.e. low values).
VGG-SS features 221 categories and IS3 130 categories which are a subset from the VGG-SS ones. However, there are many categories highly related semantically: e.g. there are 6 categories related to dog (e.g. ‘dog barking’, ‘dog bow-wow’ or ‘dog howling’), 3 related to train (‘train wheels squealing’, ‘train whistling’ and ‘train horning’), 2 related to woman (‘female singing’ and ‘female speech, woman speaking’), and so on. These categories are impossible to distinguish from each other within the visual modality when we just consider a single image. Motivated by this, we decided to define a set of broad categories for both datasets in order to select a true offscreen audio for a given category that also belongs to a completely different semantic category. The 8 broad semantic categories that we propose are the following: music/musical instruments, human voice, vehicles, devices/machines, animals, weapons, nature ambient sounds and others/miscellaneous. The categories are distinct enough (e.g. animals vs. instruments instead of cats vs. dogs) and thus offscreen sounds are easily distinguishable from onscreen sounds so that a reasonable VSSL model should be able to discriminate them.
In the case of noise, we generate a sequence of random values with normal distribution with zero mean and standard deviation 1. These values are then clipped to the range [-1, 1]. For silence, we pair the images with an empty audio. Our extended VGG-SS constitutes a total of 5,537 images, 4.61 hours of positive audio and 13.84 hours of negative audio. For the extended IS3 we have 3,242 distinct images and 6,484 audio files, corresponding to 5.40 hours of positive audio (one audio for each of the two visible sources in each image) and 16.21 hours of negative audio. For a fair comparison, we ensure that all models are tested against the same noise and offscreen audio samples. Given that the randomness in the assignments of offscreen sounds and noise generation can affect our experiments, we repeat them 10 times with different seeds and report the average metrics.
@@ II-3 F and F
| Test set | Model | cIoU U. th. | cIoU Adap. | AUC U. th. | AUC Adap. | pIA | AUC | pIA | AUC | pIA | AUC | F | F | |||
| Extended VGG-SS | RCGrad [24] | 11.71 | 37.05 | 12.07 | 37.26 | 4.42 | 95.80 | 4.43 | 95.80 | 4.41 | 95.81 | 20.86 | 20.86 | |||
| LVS [3] | 2.26 | 38.59 | 4.45 | 40.38 | 1.51 | 98.39 | 0.07 | 99.91 | 0.54 | 99.39 | 4.43 | 8.52 | ||||
| EZ-VSL [15] | 12.79 | 44.74 | 14.28 | 43.08 | 0.34 | 99.63 | 0.34 | 99.63 | 1.86 | 98.05 | 22.66 | 24.96 | ||||
| FNAC [23] | 12.74 | 48.03 | 14.26 | 44.35 | 3.81 | 96.19 | 3.81 | 96.19 | 2.70 | 97.27 | 22.50 | 24.84 | ||||
| SLAVC [16] | 18.28 | 49.64 | 19.54 | 45.90 | 0.84 | 99.16 | 0.84 | 99.16 | 4.73 | 95.23 | 30.81 | 32.57 | ||||
| SSL-TIE [12] | 18.67 | 49.86 | 19.57 | 47.05 | 0.52 | 99.42 | 0.45 | 99.52 | 1.98 | 97.93 | 31.41 | 32.68 | ||||
| SSL-Align [22] | 16.65 | 53.10 | 17.79 | 46.62 | 1.05 | 98.89 | 0.29 | 99.67 | 1.79 | 98.09 | 28.50 | 30.16 | ||||
| Extended IS3 | RCGrad [24] | 0.10 | 1.07 | 2.54 | 3.57 | 4.49 | 95.81 | 4.49 | 95.79 | 4.47 | 95.83 | 0.20 | 0.20 | |||
| LVS [3] | 1.95 | 39.38 | 4.19 | 41.07 | 1.02 | 98.85 | 0.03 | 99.96 | 0.20 | 99.77 | 3.82 | 8.04 | ||||
| EZ-VSL [15] | 3.47 | 42.08 | 5.55 | 42.66 | 0.30 | 99.67 | 0.30 | 99.67 | 0.70 | 99.22 | 6.72 | 10.51 | ||||
| FNAC [23] | 10.03 | 49.01 | 11.76 | 45.71 | 3.27 | 96.69 | 3.27 | 96.69 | 2.00 | 97.97 | 18.18 | 20.97 | ||||
| SLAVC [16] | 10.42 | 45.05 | 12.16 | 42.96 | 0.68 | 99.29 | 0.68 | 99.29 | 4.03 | 95.95 | 18.83 | 21.64 | ||||
| SSL-TIE [12] | 17.39 | 51.53 | 17.39 | 47.49 | 0.28 | 99.66 | 0.23 | 99.72 | 2.05 | 97.85 | 29.58 | 31.15 | ||||
| SSL-Align [22] | 22.54 | 63.10 | 23.51 | 52.40 | 1.86 | 97.98 | 0.45 | 99.49 | 0.82 | 99.11 | 36.72 | 37.99 | ||||
@@
As it can be observed in Fig. 1, most models present a strong overlap in the range of maximum similarities for the cases of positive, silence and offscreen audio, thus making it difficult to choose a universal threshold in which the negative audios will not be activated in the localization map. SSL-TIE and SSL-Align are the ones that present a higher discriminatory power. SSL-TIE is the most discriminative model in VGG-SS. In IS3, SSL-TIE is good at discriminating silence and noise but less good in case of offscreen. On the other hand, SSL-Align is good at discriminating noise and offscreen in IS3 but less good at silence. In light of the results in Fig. 1, we decide to set the universal localization threshold for each model as the highest 75th percentile, or 3rd quartile, of the maximum similarity values of the three types of negative audio. This value is computed in the extended test set of VGG-SS and is shown in Fig. 1 (both in numbers and as a horizontal line). We use this threshold to binarize the audio-visual similarity maps and thus obtain the localization map used in all metrics. This threshold used in the extended test sets of both VGG-SS and IS3; in this latter case with the aim of assessing if it generalizes. The threshold is computed differently for RCGrad model, as it uses gradCAM, with output values normalized in the interval . In this case, we use a threshold of .
@@ III Experiments
We conducted our experiments on the extended test sets of VGG-SS [3] and IS3 [22] datasets. Table I contains the results on the positive and each of the three negative audio inputs as well as the global metrics for both types of inputs, the best results are in bold and the second best underlined. All metrics except for two have been computed with our proposed universal threshold, U. th., (introduced in Section II). On the other hand, cIoU Adap. and AUC Adap. use the adaptive threshold proposed in [22]. In the case of positive audio input and IS3 test set, the first and second best models are SSL-Align and SSL-TIE, respectively, regardless of the threshold used to get the localization maps. In contrast, in VGG-SS, the best model with the universal threshold is SSL-TIE and SLAVC second. With the adaptive threshold, i.e. the one that uses the prior knowledge of the sounding object’s size, the best results are achieved by SSL-Align for cIoU and by SSL-TIE for AUC. EZ-VSL is the best in silence in VGG-SS and second best in IS3, while SSL-TIE is best in IS3 and second best in VGG-SS. LVS is best in noise and offscreen for both datasets. SSL-Align is second best in noise in VGG-SS and SSL-TIE in IS3. EZ-VSL is second best in offscreen in IS3 and SSL-Align in VGG-SS. Although, LVS and EZ-VSL have poor results in positive audio inputs, thus providing the worst global metrics. The global performance for both positive and negative audio is measured by F and F. We observe that best overall model in VGG-SS is SSL-TIE and in IS3 is SSL-Align, while SSL-TIE is the second best in IS3 and SLAVC in VGG-SS. The results in Table I are consistent with the analysis shown in Figure 1, where the best models according to the global metrics are the ones with less overlap in the IQR of maximum similarity values in the positive and negative cases , i.e. the more discriminatory ones.
@@
A qualitative example is shown in Figure 2. Each row represents a different model, and each column corresponds to different audio inputs paired with the same image. The example is from IS3 and there are five different positive audio files: ‘man’, ‘ukulele’ and the three types of negative audio samples. To show the effect of the universal threshold we set to zero the audio-visual similarity values below that threshold. The resulting values are then normalized to the range and they are combined with the original image. Although SSL-Align is the best overall model in IS3, it fails in this example with the audio corresponding to ‘man’ and ‘silence’. SSL-TIE correctly localizes part of the regions corresponding to both positive audio and produces empty maps for the negative audio samples. EZ-VSL, SLAVC and FNAC produce empty maps for the negatives as well as for the man. In contrast, LVS highlights almost the same region for three different audio inputs: ‘man’, ‘ukulele’ and ‘silence’. Finally, RCGrad outputs the same localization regardless of the audio input.
@@
Figure 3 shows the IoU values computed between different pairs of localization maps (positive vs. each type of negatives and negatives vs. negatives). In an effective model, IoU values should be low for positive vs. negative and high for negative vs. negative, since it should only identify sounding regions in positive cases and produce empty maps for negative ones. In VGG-SS, SSL-TIE presents the lowest IoU for positive vs. negatives (and a more pronounced difference with respect to the IoU of negatives vs. negatives). In IS3, according to these same observations, SSL-Align is the best and SSL-TIE is the second best. These results are in accordance with the F values reported in Table I.